# You Name It, I Run It: An LLM Agent to Execute Tests of Arbitrary Projects

### v1

@@ Abstract.
The ability to execute the test suite of a project is essential in many scenarios, e.g., to assess code quality and code coverage, to validate code changes made by developers or automated tools, and to ensure compatibility with dependencies. Despite its importance, executing the test suite of a project can be challenging in practice because different projects use different programming languages, software ecosystems, build systems, testing frameworks, and other tools. These challenges make it difficult to create a reliable, universal test execution method that works across different projects. This paper presents ExecutionAgent, an automated technique that installs arbitrary projects, configures them to run test cases, and produces project-specific scripts to reproduce the setup. Inspired by the way a human developer would address this task, our approach is a large language model-based agent that autonomously executes commands and interacts with the host system. The agent uses meta-prompting to gather guidelines on the latest technologies related to the given project, and it iteratively refines its process based on feedback from the previous steps. Our evaluation applies ExecutionAgent to 50 open-source projects that use 14 different programming languages and many different build and testing tools. The approach successfully executes the test suites of 33/55 projects, while matching the test results of ground truth test suite executions with a deviation of only 7.5%. These results improve over the best previously available technique by 6.6x. The costs imposed by the approach are reasonable, with an execution time of 74 minutes and LLM costs of 0.16 dollars, on average per project. We envision ExecutionAgent to serve as a valuable tool for developers, automated programming tools, and researchers that need to execute tests across a wide variety of projects.
copyright: none
@@ 1. Introduction
Executing the test suite of a software project is a critical step in various activities during software development and software engineering research. For human developers who are contributing to open-source projects, running the tests before submitting a pull request ensures that their changes do not introduce regressions. Likewise, the increasing popularity of large language model (LLM) agents that autonomously edit a project’s code (Bouzenia et al., 2024a; Zhang et al., 2024; Liu et al., 2024; Tao et al., 2024; Yang et al., 2024) creates a huge demand for a feedback mechanism to validate modifications (Spiess et al., 2025), and executing the test suite provides such a mechanism. Finally, researchers also depend on running tests, e.g., to evaluate the effectiveness of dynamic analyses (Eghbali and Pradel, 2022) or to create benchmarks involving test execution, such as Defects4J (Just et al., 2014), SWE-Bench (Jimenez et al., 2023), and DyPyBench (Bouzenia et al., 2024b).
@@
Unfortunately, executing the tests of an arbitrary project is far from straightforward in practice. Projects are developed in different programming languages and ecosystems, each with its own set of tools, dependencies, and conventions. Complex dependencies can pose significant challenges, especially when specific versions of libraries or tools are required. Documentation is often incomplete, inconsistent, or entirely missing, forcing developers to infer the necessary steps. Moreover, projects may have implicit assumptions about the environment, such as operating system specifics or required system configurations, that are not explicitly stated. The diversity of testing frameworks further complicates the process, as each framework has a unique setup and execution procedure.
@@
Currently, there are three primary methods for executing the tests of a given project, each with notable limitations. The first method involves manually following the project’s documentation and resolving any issues through trial-and-error. This approach is time-consuming and does not scale well with the number of projects. The second method is to reuse existing continuous integration/continuous deployment (CI/CD) workflows employed by project maintainers. However, not all projects have such workflows, and even when they exist, their execution often depends on a specific CI/CD platform, such as GitHub Actions, which is not fully accessible to the public. Directly using CI/CD workflows is further complicated by the fact that there are several popular platforms, each with its own configuration scripts and technology stack. The third method is to implement an automated, heuristic script designed to cover common cases. Such scripts typically focus on a single programming language and ecosystem. They lack the flexibility to handle arbitrary projects, as evidenced by studies highlighting the limitations of such scripts in diverse environments. For example, Flapy (Gruber and Fraser, 2023) works only on Python projects that are available on PyPy, and even within the Python ecosystems fails to successfully run many test suites. Another example is the pipeline used to create GitBug-Java (Silva et al., 2024), which starting from 66,042 repositories eventually managed to execute the tests of only 228 repositories.
@@
An effective solution would need to address multiple challenges. First, the approach should be aware of the latest technologies and tools for a range of popular programming languages. Second, the technique should be capable of understanding incomplete and partially outdated documentation, as often encountered in practice. Third, the approach needs a way to interact with the system, such as executing commands, monitoring outputs, and handling errors. Finally, the approach must be able to assess whether the setup process has been successful, and if not, address any problems until the test suite runs successfully. To the best of our knowledge, there is currently no existing technique that addresses these challenges. This gap presents a substantial obstacle for developers, automated coding techniques, and researchers who need a reliable and scalable solution for running tests across a wide variety of projects. Motivated by these challenges, the question addressed in this work is: How can we automatically install and run the test suite of an arbitrary project?
@@
This paper presents ExecutionAgent, the first LLM-based agent for autonomously setting up arbitrary projects and executing their test suites. The approach addresses the four challenges described above as follows. For the first challenge, we present the novel concept of meta-prompting, which asks a recently trained LLM to automatically generates up-to-date, technology-specific, and programming language-specific guidelines instead of manually engineering and hard-coding a prompt. We address the second challenge by using the LLM to parse and understand the project’s documentation and web resources related to the project. To handle the third challenge, we connect the agent to a set of tools, such as a terminal, to execute commands, monitor outputs, and interact with the system. Finally, we address the fourth challenge by enabling the agent to iteratively refine its process based on feedback from previous steps, similar to how a human developer would work.
@@
To evaluate ExecutionAgent, we apply it to 50 open-source projects that use 14 different programming languages and many different build and testing tools. The approach successfully sets up and executes the test suites of 33/50 projects. Comparing to several baselines, such as manually designed and LLM-generated scripts targeting projects in a specific language, as well as a general-purpose LLM-based agent, we find ExecutionAgent to outperform the best available technique by 6.6x. To validate the test executions, we compare them against a manually established ground truth and find that the test results (in terms of passed, failed, and skipped tests) closely resemble the ground truth, with an average deviation of 7.5%. Studying the costs of the approach, we find that the average execution time is 74 minutes per project, and the average LLM costs are 0.16 dollars per project. Overall, these results demonstrate that ExecutionAgent has the potential to serve as a valuable tool for developers, automated programming tools, and researchers that need to execute tests across a wide variety of projects.
@@
    The novel concept of meta-prompting, which allows the agent to query an LLM for the latest guidelines and technologies related to installing and running test suites.
@@ 2.1. Problem Statement
The problem we aim to address is the following: Given a software project, identified, e.g., by a URL of a git repository or a path on the local disk, we want to automatically generate scripts to setup and run the tests of the project. Specifically, the desired output consists of two scripts: one to create an isolated environment, such as a container, and one to setup up the project and run its tests within the isolated environment. By running the tests, we mean executing the project’s test suite and collecting the results, such as the number of tests that run, pass, and fail.
@@ 2.2. Overview of ExecutionAgent
We address the problem of executing the test suite of arbitrary projects using an approach that leverages LLMs and the concept of LLM agents. By LLM agent we here mean a system that uses an LLM to autonomously interact with a set of tools in order to achieve a specific goal. The tools we use are similar to those a human tasked with setting up a project might use, such as commands available via a terminal. The intuition behind this approach is that the LLM can “understand” various sources of information, such as project documentation, existing scripts, and the output of tools, and use this understanding to decide on the steps necessary to execute the tests of the given project. As shown in Figure 1, the approach encompasses two phases, which we briefly describe in the following.
@@
Given a project repository, this phase gathers information required to construct the initial prompt of the agent, called prompt ingredients. A key challenge is creating these prompt ingredients in a technology-agnostic way, as the project may be written in any programming language and use any testing framework. Our approach addresses this challenge through meta-prompting, a novel concept that allows the agent to query an LLM for the latest guidelines and technologies related to installing and running test suites. Specifically, ExecutionAgent uses meta-prompting to generate language-specific guidelines, guidelines about the latest containerization technology, and possible locations of commonly used CI/CD scripts. In addition, the approach queries the web to gather hints on installing the given project. The gathered information is then passed as prompt ingredients to the second phase.
@@
This phase repeatedly invokes tools, as guided by the LLM, in order to install the project and execute its test suite. Specifically, ExecutionAgent repeatedly iterates through three steps. Step  1 queries the LLM for the next command to execute, using a dynamically updated command prompt that contains the prompt ingredients from Phase 1 and a summary of the already invoked commands. Step  2 executes the command suggested by the LLM by invoking one of the tools. Because the output of a tool may be verbose and contain irrelevant information, step  3 requests the LLM to summarize the output and extract the most relevant information. The summarized output is then used to update the command prompt, and the three steps repeat until the agent determines that the test suite has been successfully executed. This entire process is managed by a component we call the control center.
@@
0:  Scripts to setup and run tests
@@ 2.3. Preparation Phase
The preparation phase is motivated by two goals: (1) to gather project-specific information that could be helpful for installing the given project and running its tests, and (2) to obtain guidelines that will help the LLM agent in invoking the right tools and commands during the feedback loop in phase 2. To reach these goals, we follow a set of steps that are detailed in Algorithm 1 between lines 1 and 6, and which we present in detail in the following. Since several of these steps use the idea of meta-prompting, we begin by explaining this concept.
@@ 2.3.1. Meta-Prompting
A key challenge faced by ExecutionAgent is that technologies and best practices for installing software projects and running test suites are constantly evolving. One way to address this challenge would be to spend significant amounts of time in engineering suitable prompts for the LLM agent. Such prompts should provide guidelines on how to use common build tools, package managers, testing frameworks, etc. for all the programming languages, software ecosystems, and tools that the agent might encounter. However, this approach comes with several drawbacks: it is time-consuming, is unlikely to cover all relevant technologies, and may become outdated quickly.
Instead of engineering such prompts manually, we leverage the impressive knowledge of LLMs to generate prompts with up-to-date guidelines targeted at the given project. Since LLMs are often trained on a vast range of documents, including documents describing the latest technological developments, they provide a built-in mechanism for adapting to evolving technology landscapes. We leverage this capability by using a high-level prompt, called a meta-prompt, to query the LLM for technology-specific guidelines and other information, which we then use to construct the command prompt for the agent. Specifically, ExecutionAgent uses meta-prompting to generate three kinds of prompt ingredients, which we describe in the following.
2.3.2. Language-Specific Guidelines
During initial experiments, we observed that the LLM agent benefits from guidelines that describe how to typically install and run tests of projects written in a specific programming language. Adopting the idea of meta-prompting, ExecutionAgent creates such guidelines by querying an up-to-date LLM to generate a list of language-specific guidelines. To this end, we heuristically determine the main programming language of the project (using GitHub’s Linguist library [^1] https://github.com/github-linguist/linguist) and use this information to query the LLM. Figure 2 illustrates an example of a meta-prompt for a Java project and the response from the LLM.
Guidelines Meta Prompt (for Java) I want to install and run tests of a Java project. Give a list of 10 guidelines that I should follow to achieve my task (e.g, to figure out installation steps, dependencies, how to run tests) while considering latest practices and technologies.
@@ 2.3.3. Container Guidelines
To ensure that the project’s tests are executed in an isolated environment, ExecutionAgent aims at installing and testing the project inside a container. Similar to the language-specific guidelines, we use meta-prompting to query the LLM for guidelines on the latest and most used containerization technology. For example, at the moment of writing, the GPT-4o model responds to our meta-prompt by stating that Docker is the best containerization technology for setting up an isolated environment and by providing advice on using Docker.
@@ 2.3.4. Existing CI/CD Scripts
Some projects include CI/CD scripts that provide hints about installation steps and project dependencies. To benefit from such scripts, ExecutionAgent searches for such files in the repository. While sometimes outdated and incomplete, the existing scripts may give hints about the overall process and non-standard steps to take. Instead of hard-coding specific file paths, such as .github/workflows/main.yml, ExecutionAgent uses meta-prompting to query the LLM for common locations, file names, and file extensions of CI/CD scripts within a repository. Based on the response, the approach searches for such files within the project. Because the files identified may be long and contain irrelevant information, we try to extract the most relevant parts of the scripts. To this end, ExecutionAgent queries an LLM with the raw file, asking the model to extract four kinds of information: a short natural language summary of the script, a list of dependencies of the project, any commands that are important for setting up and testing the project, and a list of relevant files or links. We call these four pieces of information a summary. Figure 3 shows an example of the LLM’s response to the summarization prompt.

### v2

@@ Abstract.
The ability to execute the test suite of a project is essential in many scenarios, e.g., to assess code quality and code coverage, to validate code changes made by developers or automated tools, and to ensure compatibility with dependencies. Despite its importance, executing the test suite of a project can be challenging in practice because different projects use different programming languages, software ecosystems, build systems, testing frameworks, and other tools. These challenges make it difficult to create a reliable, universal test execution method that works across different projects. This paper presents ExecutionAgent, an automated technique that prepares scripts for building an arbitrary project from source code and running its test cases. Inspired by the way a human developer would address this task, our approach is a large language model (LLM)-based agent that autonomously executes commands and interacts with the host system. The agent uses meta-prompting to gather guidelines on the latest technologies related to the given project, and it iteratively refines its process based on feedback from the previous steps. Our evaluation applies ExecutionAgent to 50 open-source projects that use 14 different programming languages and many different build and testing tools. The approach successfully executes the test suites of 33/50 projects, while matching the test results of ground truth test suite executions with a deviation of only 7.5%. These results improve over the best previously available technique by 6.6x. The costs imposed by the approach are reasonable, with an execution time of 74 minutes and LLM costs of USD 0.16, on average per project. We envision ExecutionAgent to serve as a valuable tool for developers, automated programming tools, and researchers that need to execute tests across a wide variety of projects.
large language models, LLM agents, autonomous software development, project setup automation, test suite execution, devops, software testing, artificial intelligence
copyright: ccdoi: 10.1145/3728922journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: ISSTAarticle: ISSTA047publicationmonth: 7ccs: Software and its engineeringccs: Software and its engineering Development frameworks and environments
@@ 1. Introduction
Executing the test suite of a software project is a critical step in various activities during software development and software engineering research. For human developers who are contributing to open-source projects, running the tests before submitting a pull request ensures that their changes do not introduce regressions. Likewise, the increasing popularity of large language model (LLM) agents that autonomously edit a project’s code (Bouzenia et al., 2024a; Zhang et al., 2024; Liu et al., 2024; Tao et al., 2024; Yang et al., 2024) creates a huge demand for a feedback mechanism to validate modifications (Spiess et al., 2025; Roychoudhury et al., 2025), and executing the test suite provides such a mechanism. Finally, researchers also depend on running tests, e.g., to evaluate the effectiveness of dynamic analyses (Eghbali and Pradel, 2022) or to create benchmarks involving test execution, such as Defects4J (Just et al., 2014), SWE-Bench (Jimenez et al., 2023), and DyPyBench (Bouzenia et al., 2024b).
@@
Unfortunately, executing the tests of an arbitrary project is far from straightforward in practice. Projects are developed in various programming languages and ecosystems, each with its own set of tools, dependencies, and conventions. Complex dependencies can pose significant challenges, especially when specific versions of libraries or tools are required. Documentation is often incomplete, inconsistent, or entirely missing, forcing developers to infer the necessary steps. Moreover, projects may have implicit assumptions about the environment, such as operating system specifics or required system configurations, that are not explicitly stated. The diversity of testing frameworks further complicates the process, as each framework has a unique setup and execution procedure.
@@
Currently, there are three primary methods for executing the tests of a given project, each with notable limitations. The first method involves manually following the project’s documentation and resolving any issues through trial-and-error. This approach is time-consuming and does not scale well with the number of projects. The second method is to reuse existing continuous integration/continuous deployment (CI/CD) workflows employed by project maintainers. However, not all projects have such workflows, and even when they exist, their execution often depends on a specific CI/CD platform, such as GitHub Actions, which is not fully accessible to the public. Directly using CI/CD workflows is further complicated by the fact that there are several popular platforms, each with its own configuration scripts and technology stack. The third method is to implement an automated, heuristic script designed to cover common cases. Such scripts typically focus on a single programming language and ecosystem, and they lack the flexibility to handle arbitrary projects: For example, Flapy (Gruber and Fraser, 2023) works only on Python projects that are available on PyPy, and even within the Python ecosystems fails to successfully run many test suites. Another example is the pipeline used to create GitBug-Java (Silva et al., 2024), which, starting from 66,042 repositories, eventually managed to execute the tests of only 228 repositories.
@@
An effective solution would need to address multiple challenges. First, the approach should be aware of the latest technologies and tools for a range of popular programming languages. Second, the approach should be capable of understanding incomplete and partially outdated documentation. Third, the approach needs a way to interact with the system, such as executing commands, monitoring outputs, and handling errors. Finally, the approach must be able to assess whether the setup process has been successful, and if not, address any problems until the test suite runs successfully. To the best of our knowledge, there is currently no existing work that addresses these challenges. This gap presents a substantial obstacle for developers, automated coding techniques, and researchers who need a reliable and scalable solution for running tests across a wide variety of projects. Motivated by these challenges, the question addressed in this work is: Given an arbitrary project, how can we automatically build the project and run its test suite?
@@
This paper presents ExecutionAgent, the first LLM-based agent for autonomously setting up arbitrary projects and executing their test suites. The approach addresses the four challenges described above as follows. For the first challenge, we present the novel concept of meta-prompting, which asks a recently trained LLM to automatically generate up-to-date, technology-specific, and programming language-specific guidelines instead of manually engineering and hard-coding a prompt. We address the second challenge by using the LLM to parse and understand the project’s documentation and web resources related to the project. To handle the third challenge, we connect the agent to a set of tools, such as a terminal, to execute commands, monitor outputs, and interact with the system. Finally, we address the fourth challenge by enabling the agent to iteratively refine its process based on feedback from previous steps, similar to how a human developer would work.
@@
To evaluate ExecutionAgent, we apply it to 50 open-source projects that use 14 different programming languages and many different build and testing tools. The approach successfully sets up and executes the test suites of 33/50 projects. Comparing to several baselines, such as manually designed and LLM-generated scripts targeting projects in a specific language, as well as a general-purpose LLM-based agent, we find ExecutionAgent to outperform the best available technique by 6.6x. To validate the test executions, we compare them against a manually established ground truth and find that the test results (in terms of the number of passed, failed, and skipped tests) closely resemble the ground truth, with an average deviation of 7.5%. Studying the costs of the approach, we find that the average execution time is 74 minutes per project, and the average LLM costs are USD 0.16 per project. Overall, these results demonstrate that ExecutionAgent has the potential to serve as a valuable tool for developers, automated programming tools, and researchers that need to execute tests across a wide variety of projects.
@@
    The novel concept of meta-prompting, which allows the agent to query an LLM for the latest guidelines and technologies related to building projects and running test suites.
@@ 2.1. Problem Statement
The problem we aim to address is the following: Given a software project, identified, e.g., by a URL of a git repository or a file path, we want to automatically generate scripts to install the project and run its tests. Specifically, the desired output consists of two scripts: one to create an isolated environment, such as a container, and one to build the project and run its tests within the isolated environment. By running the tests, we mean executing the project’s test suite and collecting the results, such as the number of tests that pass, fail, and are skipped.
@@ 2.2. Overview of ExecutionAgent
We address the above problem by presenting an approach that leverages LLMs and the concept of LLM agents. By LLM agent we here mean a system that uses an LLM to autonomously interact with a set of tools in order to achieve a specific goal (Roychoudhury et al., 2025). The tools we use are similar to those a human tasked with setting up a project might use, such as commands available via a terminal. The intuition behind this approach is that the LLM can “understand” various sources of information, such as project documentation, existing scripts, and the output of tools, and use this understanding to decide on the steps necessary to execute the tests of the given project. As shown in Figure 1, the approach encompasses two phases, which we briefly describe in the following.
@@
Given a project repository, this phase gathers information required to construct the initial prompt of the agent, called prompt ingredients. A key challenge is creating these prompt ingredients in a technology-agnostic way, as the project may be written in any programming language and use any testing framework. Our approach addresses this challenge through meta-prompting, a novel concept that allows the agent to query an LLM for the latest guidelines and technologies related to building projects and running test suites. Specifically, ExecutionAgent uses meta-prompting to generate language-specific guidelines, guidelines about the latest containerization technology, and possible locations of commonly used CI/CD scripts. In addition, the approach queries the web to gather hints on building the given project. Finally, our approach uses an LLM to generate general-purpose, language-independent guidelines for creating a build and test environment for an arbitrary project. These guidelines help the agent follow a high-level plan to reach its goals. The generated and gathered information is then passed as prompt ingredients to the second phase.
@@
This phase repeatedly invokes tools, as guided by the LLM, in order to build the project and execute its test suite. Specifically, ExecutionAgent repeatedly iterates through three steps. Step  1 queries the LLM for the next command to execute, using a dynamically updated command prompt that contains the prompt ingredients from Phase 1 and a summary of the already invoked commands. Step  2 executes the command suggested by the LLM by invoking one of the tools. Because the output of a tool may be verbose and contain irrelevant information, step  3 requests the LLM to summarize the output and extract the most relevant information. The summarized output is then used to update the command prompt, and the three steps repeat until the agent determines that the test suite has been successfully executed. This entire process is managed by a component we call the control center.
@@
0:  Scripts to set up an environment and run tests
@@ 2.3. Preparation Phase
The preparation phase is motivated by two goals: (1) to gather project-specific information that could be helpful for building the given project and running its tests, and (2) to obtain guidelines that will help the LLM agent invoke the right tools and commands during the feedback loop in phase 2. To reach these goals, we follow a set of steps that are detailed in Algorithm 1 between lines 1 and 6, and which we present in detail in the following. Since several of these steps use the idea of meta-prompting, we begin by explaining this concept.
@@ 2.3.1. Meta-Prompting
A key challenge faced by ExecutionAgent is that technologies and best practices for building software projects and running test suites are constantly evolving. One way to address this challenge would be to spend significant amounts of time in engineering suitable prompts for the LLM agent. Such prompts should provide guidelines on how to use common build tools, package managers, testing frameworks, etc. for all the programming languages, software ecosystems, and tools that the agent might encounter. However, this approach comes with several drawbacks: it is time-consuming, unlikely to cover all relevant technologies, and may become outdated quickly.
Instead of engineering such prompts manually, we leverage the impressive knowledge of LLMs to generate prompts with up-to-date guidelines targeted at the given project. Since LLMs are often trained on a vast range of documents, including documents describing the latest technological developments, they offer a built-in mechanism for adapting to evolving technology landscapes. We leverage this capability by using a high-level prompt, called a meta-prompt, to query the LLM for technology-specific guidelines and other information, which we then use to construct the command prompt for the agent. Specifically, ExecutionAgent uses meta-prompting to generate three kinds of prompt ingredients, which we describe in the following.
2.3.2. Language-Specific and Language-Independent Guidelines
During initial experiments, we observed that the LLM agent benefits from guidelines that describe how to typically build projects written in a specific language and how to run their tests. Adopting the idea of meta-prompting, ExecutionAgent creates such guidelines by querying an up-to-date LLM to generate a list of language-specific guidelines. To this end, we heuristically infer the main programming language of the project during the preparation phase (using GitHub’s Linguist library [^1] https://github.com/github-linguist/linguist) and use this information to generate the language-specific guidelines.
Guidelines Meta Prompt (for Java) I want to install and run tests of a Java project. Give a list of 10 guidelines that I should follow to achieve my task (e.g., to figure out installation steps, dependencies, how to run tests) while considering latest practices and technologies.
@@ 2.3.2. Language-Specific and Language-Independent Guidelines
Figure 2 illustrates an example of a meta-prompt for a Java project and the response from the LLM. The prompt asks for a list of ten guidelines for a specific programming language. The reason for asking for a specific number of guidelines is that otherwise, the LLM sometimes generates only a few broad guidelines, which are lacking in detail, or may generate an extensively long list, which is hard to process. To avoid such variation, we explicitly request a specific number, and we found ten to be sufficiently comprehensive.
In addition to creating language-specific guidelines, we instruct the agent to collect further information about the projects and tests using a set of language-independent guidelines that we also generate with an LLM. These language-independent guidelines ensure that our approach remains effective even when a project’s setup does not strictly adhere to the common practices of its main language, for instance, due to significant influence from secondary languages.
@@ 2.3.3. Container Guidelines
To ensure that the project’s tests are executed in an isolated environment, ExecutionAgent aims at building and testing the project inside a container. Similar to the language-specific guidelines, we use meta-prompting to query the LLM for guidelines on the latest and most used containerization technology. For example, at the moment of writing, the GPT-4o model responds to our meta-prompt by stating that Docker is the best option.
@@ 2.3.4. Existing CI/CD Scripts
Some projects include CI/CD scripts that provide hints about installation steps and project dependencies. To leverage such scripts, ExecutionAgent follows a three-step process. First, it uses meta-prompting to query the LLM for common folder names, file names, and extensions associated with CI/CD scripts. Second, based on the obtained list, it searches the repository for matching files and folders. Third, the identified files are analyzed by the LLM to determine which files are relevant to building and testing.
While CI/CD scripts can sometimes be outdated or incomplete, they may still provide useful insights into the overall setup process and any required non-standard steps. To extract the most relevant information, ExecutionAgent queries an LLM with the raw file contents, asking it to provide: (1) a concise natural language summary of the script, (2) a list of project dependencies, (3) important commands for setting up and testing, and (4) references to relevant files or links. Figure 3 shows an example of the LLM’s response to the summarization prompt.